# On Statistical Inference for High-Dimensional Binary Time Series

### v1

@@ 1 Introduction
Time series data frequently appear in practice, making their analysis a common and essential task in many real-life applications, like those introduced in Lütkepohl (2005), Politis and McElroy (2019), Berry and West (2020), Das and Politis (2021), Dette and Wu (2022), Yuan and Shou (2024), Ma et al. (2024), among others. After practitioners collect a time series dataset, various models, such as ARMA or GARCH processes, are often resorted to summarize the conditional mean or volatility information of the data. Furthermore, after fitting a suitable model, statisticians may perform statistical inference on the underlying data generating process, like those in Section 15.4 of Lütkepohl (2005), or perform mean or probabilistic forecasting, like those in Zhang et al. (2025).
@@
Despite wide applications and advantages of the existing models, an issue with these models is that they are mainly designed for time series with real value support—making them unsuitable for time series data that exhibit special characteristics. For example, if the time series dataset has binary support, fitting an autoregressive model does not guarantee that maintain such characteristics. However, constraints on the support of practical time series datasets often exist, which may arise from the physical meaning of the data. For example, Keenan (1982) mentioned that binary-value time series can arise when statisticians consider the dynamic occurrence of events or changes in states; Livsey et al. (2018) introduced a count time series model for hurricane data; other research like Xu et al. (2012), Fokianos et al. (2009), Fokianos and Tjøstheim (2012), and the reference therein, also raised similar issues.
@@
This manuscript leverages the work of Jentsch and Reichmann (2022) to introduce a high-dimensional generalized binary vector autoregressive (gbVAR) process for high-dimensional binary-valued vector time series data. After incorporating the sparsity assumption of the parameter matrices, it introduces a post-selection estimator to estimate the parameter matrices of the model and derives the associated Gaussian approximation for the proposed estimator. The complex spatial and temporal dependence of the vector time series data causes the proposed estimator to exhibit complex and hard-to-estimate variances and covariances, making it unrealistic to perform statistical inference through manual calculations. Concerning this, we propose a bootstrap algorithm, named the second-order wild bootstrap, to assist statisticians in performing statistical inference on the parameter matrices. This manuscript also theoretically justifies the validity of the proposed bootstrap algorithm.
@@
It is common for modern time series to simultaneously exhibit both high dimensionality and constraints on their support. Given the prevalence of non-real-valued-supported data—such as binary vector time series—in various application domains, including Kharin et al. (2018); Darolles et al. (2019); Franchi and Truquet (2024), our work is expected to be of interest to practitioners in these domains.
@@
This paper is organized as follows: Section 2 makes a thorough review of the analysis of non-real-value time series data. Section 3 introduces gbVAR Models and their stochastic probability for low-dimensional structures. Section 4 proposes the transition matrix estimation procedure and bootstrap inference for binary time series. Section 5 presents asymptotic analysis based on the functional dependence measure and Gaussian approximation. Section 6 provides numerical experiments, and Section 7 summarizes discussions and conclusions. Technical proofs are deferred to the online supplement Dai and Zhang (2025).
@@ 2 Related literature
Analysis of non-real-value-supported time series. There have been numerous attempts in the literature to model time series whose support is not the real axis. To mention a few, the work of McKenzie (1985); Al-Osh and Alzaid (1987) introduced the integer-value autoregressive processes, which replaced the conventional scalar multiplication with the so-called “thinning operators,” to accommodate integer-value time series. McKenzie (1988); Al-Osh and Alzaid (1988) later extended this approach to develop integer-value moving average processes. More recent studies on integer-value autoregressive processes include Nastić et al. (2016); Zhang et al. (2023a); Ristić et al. (2016); Huh et al. (2017); Chen et al. (2020), which introduced random coefficients, new innovation structures, and models incorporating conditional heteroskedasticity. In addition to the aforementioned work, we refer readers to Bourguignon and Weiß (2017); Weiß (2021); Fokianos et al. (2022); Chen et al. (2022); Piancastelli et al. (2023), for further exploration of this topic.
@@
Other attempts leveraged generalized linear models, such as the Logistic regression model, to accommodate integer-value time series. Examples include Liang and Zeger (1989); Benjamin et al. (2003) that developed logistic regression models for multivariate binary time series, Fokianos et al. (2009); Christou and Fokianos (2014); Ahmad and Francq (2016) introduced the quasi-likelihood inference for generalized count autoregressive models, while Fokianos and Kedem (2004) proposed a partial likelihood inference method for modeling the generalized time series. Zhu et al. (2017) introduced a neural autoregressive model, which was later extended to count-valued time series by Armillotta and Fokianos (2024). Other nonlinear models for count time series have been studied in Fokianos and Tjøstheim (2011); Fokianos (2012); Christou and Fokianos (2014); Wang et al. (2014); Dunsmuir (2015); Davis and Liu (2016). More recent and applied works can be found in the literature, such as Hall et al. (2018); Amillotta et al. (2022); Jia et al. (2023); Pumi et al. (2024).
@@
Analysis of high-dimensional time series. Modern-era data often exhibit complex structures. In particular, due to the advancement in data collection techniques, the dimension of the observations can be comparable to, or exceed, the number of observations. The literature has extensively explored high-dimensional real-valued time series, as seen in works by Jentsch and Politis (2015); Zhang and Wu (2017, 2021); Krampe et al. (2023); Zhou and Dette (2023), among others. Compared to the rich literature in high-dimensional real-valued time series, to the best of the authors’ knowledge, relatively little research has been conducted on high-dimensional time series with non-real number support. Recent examples include Fokianos et al. (2020); Guo and Qiao (2023); Düker et al. (2023), and the references therein.
@@
Applications to network data. One potential application of our study lies in the analysis of dynamic networks. Practitioners often assign binary random variables to the edges of network data when performing statistical inference. Classical approaches assume independence among these random variables, thereby ignoring the potential spatial and temporal dependence across edges. Recent studies have tended to solve such a problem. For example, Jiang et al. (2023) introduced a first-order autoregressive dynamic network process to introduce temporal dependence, Hanneke et al. (2010); Krivitsky and Handcock (2014); Leifeld et al. (2018) developed temporal exponential family random-graph models, Süveges and Olhede (2023); Chang et al. (2024b) respectively proposed autoregressive network models based on logistic regressions and Markov chains. We also refer readers to Mantziou et al. (2023); Zhu and Yao (2025), and references therein.
@@
This section leverages the work of Jentsch and Reichmann (2022) to introduce the gbVAR model for high-dimensional binary-valued vector time series. In addition, it explores several statistical properties of this model. Given a $d\times(p+1)d$ parameter matrix for $1\leq q\leq p$,
@@
is a diagonal matrix with positive definite $\mathcal{B}$, following Jentsch and Reichmann (2022), we define the “counterpart matrix” $\mathcal{P}_{|\cdot|}$ as follows:
@@
The remainder part of this section focuses on statistical properties of a lag-1 gbVAR model. While Jentsch and Reichmann (2019, 2022) addressed the estimation of autoregressive coefficient matrices for binary vector autoregressive processes, they did not consider cases where the dimension $d$ is moderately or very large, and did not perform statistical inference. Our work complements their studies by extending the analysis to high-dimensional settings and incorporating sparse structures in the statistical inference. Additionally, we establish simultaneous confidence intervals for the entire autoregressive coefficient matrices.
@@ 3.1 Stochastic properties of gbVAR(1) process
Proposition 3 below presents a result similar to the moving-averaging representation commonly seen in real-valued vector autoregressive models, as introduced in Brockwell and Davis (2006). The MA representation of (4) can be established under the following condition:
@@
A common way to build up model identification procedure in vector time series analysis literature involves the Yule-Walker equation, as introduced in Brockwell and Davis (1991). Proposition 5 leverages similar idea to identify $\mathcal{A}$.
@@ 4.1 Estimating Transition Matrix in VAR Models
The key difficulty of the estimation lies in the high-dimensionality of data, where the dimension $d$ may exceed the number of observations $n$. In setups where $d$ is substantially smaller than $n$, according to the Yule–Walker equation in Proposition 5, practitioners may replace the population covariance matrices $\bm{\Sigma}^{(k)},k=0,1$ with the sample covariance matrices $\widehat{\bm{\Sigma}}^{(k)},$ and obtain the estimator by inverting $\widehat{\bm{\Sigma}}^{(0)}.$ However, when $d>n,$ the sample covariance matrix $\widehat{\bm{\Sigma}}^{(0)}$ is not invertible, and the solution of the Yule-Walker equations is not well-defined. The literature has offered methods to estimate precision matrices for high-dimensional data, including the works of Cai et al. (2011); Fan et al. (2016); Zhang (2021a), among others. However, such methods often impose sparse structures on the precision matrices of data, which are not compatible with our settings, where the parameter matrix $\mathcal{A}$ is already assumed to be sparse.
@@ 4.2 Second-order wild bootstrap
Bootstrap method has emerged as a powerful and versatile tool for estimating distribution information of sample statistics, especially when identifying a pivotal quantity is difficult. Efron (1979); Wu (1986); Stine (1985), among others, introduced various bootstrap algorithms for independent data, while later works, including Politis and Romano (1992, 1994), claimed the effectiveness of bootstrap algorithms in time series analysis. Examples of recent studies on bootstrap algorithms and their extensions include (Bühlmann, 1997; Jentsch and Politis, 2015) for linear processes, (Meyer et al., 2020) for spectral statistics, (Kreiss and Paparoditis, 2015) for locally stationary time series. Among them, Shao (2010) introduced the dependent wild bootstrap for stationary time series, which can be extended to different setups, even non-stationary time series, as demonstrated in (Zhang and Politis, 2022; Zhang et al., 2023b; Zhang, 2023).
gbVAR processes demonstrate complicated spatial and temporal correlations. Compared to the vector autoregressive process, the parameter matrices in a gbVAR process are random variables instead of numbers, so the covariance matrix of the post-selection estimator depends not only on the innovations’ covariance matrix but also on the covariance structure induced by the random coefficients. These two sources of covariances inevitably lead to more complex and harder to calculate variance and covariance structure in the estimators—according to (Krampe et al., 2023), even when the innovations are i.i.d, the autoregressive coefficient estimator of gbVAR processes still has complicated covariance matrices. Our work leverages the second-order wild bootstrap introduced in Zhang (2023) to develop a bootstrap algorithm suitable for analyzing estimators of the gbVAR time series.
@@ 5 Asymptotic analysis of the estimator
Analysis of time series data typically relies on short-range dependent conditions, such as the mixing conditions in Doukhan (1994). However, the strong mixing conditions are not well-suited for integer-value time series—as illustrated in Wu (2005), even a simple Bernoulli shift process can fail to satisfy the mixing condition. Wu (2005) introduced another condition to describe short range dependence, named the “physical dependence measure,” which were later used in the work of Rho and Shao (2015), Zhang and Wu (2017), Zhang and Wu (2018), Li et al. (2023), among others. This section demonstrates that the data generating process (4) obeys the physical dependence condition and establishes the asymptotic properties of the post-selection estimator. Let $\bm{\varepsilon}_{t},$ $t\in\mathbb{Z}$ be i.i.d. random variables, define $\mathcal{F}_{-\infty}^{t}=(\ldots,\bm{\varepsilon_{t-1}},\bm{\varepsilon_{t}}).$ We suppose $X_{t}\in\mathbb{R}^{d},t\in\mathbb{Z}$ be a stationary process that satisfies equation (4). Furthermore, we assume that each $X_{t}$ is a function of $\mathcal{F}_{-\infty}^{t}$ obeying the following form
@@
The dependence parameter $\rho$ in equation (15) is restricted to $(0,1)$ to ensure the weak stationarity of DAM. If $\rho>1,$ the variance of process $X_{t}$ may diverge to infinity and measures, such as Phillips and Perron (1988) and DeJong et al. (1992) proposed tests for detecting the presence of a unit root in general time series models. Zhang and Cheng (2014) considered the special case which imposes the stronger GMC condition $\max_{1\leq s\leq d}\sum_{t=m}^{\infty}\delta_{t,q,s}\leq C\rho^{m}$ with some constant and $\rho\in(0,1).$
In Zhang and Wu (2017), the condition can be extended to exponential-type tail, such as sub-Gaussian (exponential) innovation. We can also define an adapted Orlicz norm if there exists a positive constant $\nu>0$ such that
@@ 5.1 Bernstein-type Inequality under Dependence
In this section, we introduce a univariate Bernstein-type inequality for the process in (14) with $d=1$, the assumption of boundedness and finite second DAM. The well-known Bernstein inequality (Bernstein (1946)) provides an exponential concentration result for sums of uniformly bounded independent random variables. Let $X_{1},\ldots,X_{n}$ be independent random variables with $\mathbb{E}X_{i}=0$ and $\sigma_{i}^{2}=\mathrm{Var}(X_{i})<\infty$, and $|X_{i}|\leq M$, for all $i$. Denote $S_{n}=\sum_{i=1}^{n}X_{i}$. Then, for any $x>0$, we have
@@
Proposition 14 (Theorem 2.1 in Zhang (2021b)).
@@
If $\|X_{\cdot}\|_{2}=O(1)$, compared with the classical Bernstein inequality (16) for independent processes, the result provides an additional $(\log n)^{2}$ order in the sub-exponential-type tail. Theorem 6 in Adamczak (2008) provides a slightly sharper inequality involving only an additional $\log n$ order:
@@
where $S_{n}=\sum_{i=1}^{n}X_{i}$, $X_{i}=\sum_{i=1}^{n}f(Y_{i})$, $(Y_{i})$ is a Markov chain satisfying some minorization condition, $f$ is a bounded function, and $\nu^{2}=\lim_{n\rightarrow\infty}\mathrm{Var}(S_{n}/\sqrt{n})$. The result is as sharp as Theorem 2 in Merlevede et al. (2009) up to a multiplicative constant in the exponential function:

### v2

@@ 1 Introduction
Time series data frequently appear in practice, making their analysis a common and essential task in many real-life applications, like those introduced in MR2172368, politis2019time, MR4154894, MR4270034, MR4356579, 10.1371/journal.pbio.3002758, MR4698614, among others. After practitioners collect a time series dataset, various models, such as ARMA or GARCH processes, are often resorted to summarize the conditional mean or volatility information of the data. Furthermore, after fitting a suitable model, statisticians may perform statistical inference on the underlying data generating process, like those in Section 15.4 of MR2172368, or perform mean or probabilistic forecasting, like those in ZHANG2025112611.
@@
Despite wide applications and advantages of the existing models, an issue with these models is that they are mainly designed for time series with real value support—making them unsuitable for time series data that exhibit special characteristics. For example, if the time series dataset has binary support, fitting an autoregressive model does not guarantee that maintain such characteristics. However, constraints on the support of practical time series datasets often exist, which may arise from the physical meaning of the data. For example, keenan1982time mentioned that binary-value time series can arise when statisticians consider the dynamic occurrence of events or changes in states; MR3773399 introduced a count time series model for hurricane data; other research like MR2957867, fokianos2009poisson, MR2981620, and the reference therein, also raised similar issues.
@@
This manuscript leverages the work of jentsch2022generalized to introduce a high-dimensional generalized binary vector autoregressive (gbVAR) process for high-dimensional binary-valued vector time series data. After incorporating the sparsity assumption of the parameter matrices, it introduces a post-selection estimator to estimate the parameter matrices of the model and derives the associated Gaussian approximation for the proposed estimator. The complex spatial and temporal dependence of the vector time series data causes the proposed estimator to exhibit complex and hard-to-estimate variances and covariances, making it unrealistic to perform statistical inference through manual calculations. Concerning this, we propose a bootstrap algorithm, named the second-order wild bootstrap, to assist statisticians in performing statistical inference on the parameter matrices. This manuscript also theoretically justifies the validity of the proposed bootstrap algorithm.
@@
It is common for modern time series to simultaneously exhibit both high dimensionality and constraints on their support. Given the prevalence of non-real-valued-supported data—such as binary vector time series—in various application domains, including kharin2018statistical; darolles2019bivariate; franchi2024theory, our work is expected to be of interest to practitioners in these domains.
@@
This paper is organized as follows: Section 2 makes a thorough review of the analysis of non-real-value time series data. Section 3 introduces gbVAR Models and their stochastic probability for low-dimensional structures. Section 4 proposes the transition matrix estimation procedure and bootstrap inference for binary time series. Section 5 presents asymptotic analysis based on the functional dependence measure and Gaussian approximation. Section 6 provides numerical experiments, and Section 7 summarizes discussions and conclusions. Technical proofs are deferred to the online supplement Supplement.
@@ 2 Related literature
Analysis of non-real-value-supported time series. There have been numerous attempts in the literature to model time series whose support is not the real axis. To mention a few, the work of mckenzie1985some; al1987first introduced the integer-value autoregressive processes, which replaced the conventional scalar multiplication with the so-called “thinning operators,” to accommodate integer-value time series. mckenzie1988some; al1988integer later extended this approach to develop integer-value moving average processes. More recent studies on integer-value autoregressive processes include nastic2016random; zhang2023new; ristic2016binomial; huh2017monitoring; chen2020two, which introduced random coefficients, new innovation structures, and models incorporating conditional heteroskedasticity. In addition to the aforementioned work, we refer readers to bourguignon2017inar; weiss2021stationary; fokianos2022statistical; chen2022new; piancastelli2023flexible, for further exploration of this topic.
@@
Other attempts leveraged generalized linear models, such as the Logistic regression model, to accommodate integer-value time series. Examples include liang1989class; MR1965687 that developed logistic regression models for multivariate binary time series, fokianos2009poisson; christou2014quasi; ahmad2016poisson introduced the quasi-likelihood inference for generalized count autoregressive models, while fokianos2004partial proposed a partial likelihood inference method for modeling the generalized time series. zhu2017network introduced a neural autoregressive model, which was later extended to count-valued time series by armillotta2024count. Other nonlinear models for count time series have been studied in fokianos2011log; fokianos2012count; christou2014quasi; wang2014self; dunsmuir2015generalized; davis2016theory. More recent and applied works can be found in the literature, such as hall2018learning; amillotta2022generalized; jia2023latent; pumi2024unit.
@@
Analysis of high-dimensional time series. Modern-era data often exhibit complex structures. In particular, due to the advancement in data collection techniques, the dimension of the observations can be comparable to, or exceed, the number of observations. The literature has extensively explored high-dimensional real-valued time series, as seen in works by jentsch2015covariance; zhang2017gaussian; MR4206676; krampe2023structural; MR4718536, among others. Compared to the rich literature in high-dimensional real-valued time series, to the best of the authors’ knowledge, relatively little research has been conducted on high-dimensional time series with non-real number support. Recent examples include fokianos2020multivariate; guo2023consistency; duker2023high, and the references therein.
@@
Applications to network data. One potential application of our study lies in the analysis of dynamic networks. Practitioners often assign binary random variables to the edges of network data when performing statistical inference. Classical approaches assume independence among these random variables, thereby ignoring the potential spatial and temporal dependence across edges. Recent studies have tended to solve such a problem. For example, jiang2023autoregressive introduced a first-order autoregressive dynamic network process to introduce temporal dependence, hanneke2010; krivitsky2014; Leifeld2018 developed temporal exponential family random-graph models, suveges2023; chang2024autoregressive respectively proposed autoregressive network models based on logistic regressions and Markov chains. We also refer readers to mantziou2023gnar; zhu2025autoregressive, and references therein.
@@
This section leverages the work of jentsch2022generalized to introduce the gbVAR model for high-dimensional binary-valued vector time series. In addition, it explores several statistical properties of this model. Given a $d\times(p+1)d$ parameter matrix for $1\leq q\leq p$,
@@
is a diagonal matrix with positive definite $\mathcal{B}$, following jentsch2022generalized, we define the “counterpart matrix” $\mathcal{P}_{|\cdot|}$ as follows:
@@
The remainder part of this section focuses on statistical properties of a lag-1 gbVAR model. While jentsch2019generalized; jentsch2022generalized addressed the estimation of autoregressive coefficient matrices for binary vector autoregressive processes, they did not consider cases where the dimension $d$ is moderately or very large, and did not perform statistical inference. Our work complements their studies by extending the analysis to high-dimensional settings and incorporating sparse structures in the statistical inference. Additionally, we establish simultaneous confidence intervals for the entire autoregressive coefficient matrices.
@@ 3.1 Stochastic properties of gbVAR(1) process
Proposition 3 below presents a result similar to the moving-averaging representation commonly seen in real-valued vector autoregressive models, as introduced in MR2839251. The MA representation of (4) can be established under the following condition:
@@
A common way to build up model identification procedure in vector time series analysis literature involves the Yule-Walker equation, as introduced in brockwell1991time. Proposition 5 leverages similar idea to identify $\mathcal{A}$.
@@ 4.1 Estimating Transition Matrix in VAR Models
The key difficulty of the estimation lies in the high-dimensionality of data, where the dimension $d$ may exceed the number of observations $n$. In setups where $d$ is substantially smaller than $n$, according to the Yule–Walker equation in Proposition 5, practitioners may replace the population covariance matrices $\bm{\Sigma}^{(k)},k=0,1$ with the sample covariance matrices $\widehat{\bm{\Sigma}}^{(k)},$ and obtain the estimator by inverting $\widehat{\bm{\Sigma}}^{(0)}.$ However, when $d>n,$ the sample covariance matrix $\widehat{\bm{\Sigma}}^{(0)}$ is not invertible, and the solution of the Yule-Walker equations is not well-defined. The literature has offered methods to estimate precision matrices for high-dimensional data, including the works of cai2011constrained; 10.1111/ectj.12061; 29a36494-561a-3a70-b30a-b00da391c084, among others. However, such methods often impose sparse structures on the precision matrices of data, which are not compatible with our settings, where the parameter matrix $\mathcal{A}$ is already assumed to be sparse.
@@ 4.2 Second-order wild bootstrap
Bootstrap method has emerged as a powerful and versatile tool for estimating distribution information of sample statistics, especially when identifying a pivotal quantity is difficult. efron1979bootstrap; wu1986jackknife; stine1985bootstrap, among others, introduced various bootstrap algorithms for independent data, while later works, including politis1990circular; politis1994stationary, claimed the effectiveness of bootstrap algorithms in time series analysis. Examples of recent studies on bootstrap algorithms and their extensions include (buhlmann197Sieve; jentsch2015covariance) for linear processes, (meyer2020extending) for spectral statistics, (kreiss2015bootstrapping) for locally stationary time series. Among them, shao2010dependent introduced the dependent wild bootstrap for stationary time series, which can be extended to different setups, even non-stationary time series, as demonstrated in (zhang2022ridge; zhang2023simultaneous; zhang2023statistical).
gbVAR processes demonstrate complicated spatial and temporal correlations. Compared to the vector autoregressive process, the parameter matrices in a gbVAR process are random variables instead of numbers, so the covariance matrix of the post-selection estimator depends not only on the innovations’ covariance matrix but also on the covariance structure induced by the random coefficients. These two sources of covariances inevitably lead to more complex and harder to calculate variance and covariance structure in the estimators—according to (krampe2023structural), even when the innovations are i.i.d, the autoregressive coefficient estimator of gbVAR processes still has complicated covariance matrices. Our work leverages the second-order wild bootstrap introduced in zhang2023statistical to develop a bootstrap algorithm suitable for analyzing estimators of the gbVAR time series.
@@ 5 Asymptotic analysis of the estimator
Analysis of time series data typically relies on short-range dependent conditions, such as the mixing conditions in MR1312160. However, the strong mixing conditions are not well-suited for integer-value time series—as illustrated in wu2005nonlinear, even a simple Bernoulli shift process can fail to satisfy the mixing condition. wu2005nonlinear introduced another condition to describe short range dependence, named the “physical dependence measure,” which were later used in the work of Rho03072015, zhang2017gaussian, zhang2018asymptotic, li2023inference, among others. This section demonstrates that the data generating process (4) obeys the physical dependence condition and establishes the asymptotic properties of the post-selection estimator. Let $\bm{\varepsilon}_{t},$ $t\in\mathbb{Z}$ be i.i.d. random variables, define $\mathcal{F}_{-\infty}^{t}=(\ldots,\bm{\varepsilon_{t-1}},\bm{\varepsilon_{t}}).$ We suppose $X_{t}\in\mathbb{R}^{d},t\in\mathbb{Z}$ be a stationary process that satisfies equation (4). Furthermore, we assume that each $X_{t}$ is a function of $\mathcal{F}_{-\infty}^{t}$ obeying the following form
@@
The dependence parameter $\rho$ in equation (15) is restricted to $(0,1)$ to ensure the weak stationarity of DAM. If $\rho>1,$ the variance of process $X_{t}$ may diverge to infinity and measures, such as phillips1988testing and dejong1992power proposed tests for detecting the presence of a unit root in general time series models. zhang2014bootstrapping considered the special case which imposes the stronger GMC condition $\max_{1\leq s\leq d}\sum_{t=m}^{\infty}\delta_{t,q,s}\leq C\rho^{m}$ with some constant and $\rho\in(0,1).$
In zhang2017gaussian, the condition can be extended to exponential-type tail, such as sub-Gaussian (exponential) innovation. We can also define an adapted Orlicz norm if there exists a positive constant $\nu>0$ such that
@@ 5.1 Bernstein-type Inequality under Dependence
In this section, we introduce a univariate Bernstein-type inequality for the process in (14) with $d=1$, the assumption of boundedness and finite second DAM. The well-known Bernstein inequality (bernstein1946theory) provides an exponential concentration result for sums of uniformly bounded independent random variables. Let $X_{1},\ldots,X_{n}$ be independent random variables with $\mathbb{E}X_{i}=0$ and $\sigma_{i}^{2}=\mathrm{Var}(X_{i})<\infty$, and $|X_{i}|\leq M$, for all $i$. Denote $S_{n}=\sum_{i=1}^{n}X_{i}$. Then, for any $x>0$, we have
@@
Proposition 14 (Theorem 2.1 in zhang2021robust).
@@
If $\|X_{\cdot}\|_{2}=O(1)$, compared with the classical Bernstein inequality (16) for independent processes, the result provides an additional $(\log n)^{2}$ order in the sub-exponential-type tail. Theorem 6 in adamczak2008tail provides a slightly sharper inequality involving only an additional $\log n$ order:
@@
where $S_{n}=\sum_{i=1}^{n}X_{i}$, $X_{i}=\sum_{i=1}^{n}f(Y_{i})$, $(Y_{i})$ is a Markov chain satisfying some minorization condition, $f$ is a bounded function, and $\nu^{2}=\lim_{n\rightarrow\infty}\mathrm{Var}(S_{n}/\sqrt{n})$. The result is as sharp as Theorem 2 in merlevede2009bernstein up to a multiplicative constant in the exponential function: